OpenStudy (anonymous):

A(-2,5) with a slope of 3/4 I am trying to write the general form of the line that passes through point A. I am using y-5=3/4(x+2)

OpenStudy (anonymous):

I am getting \[-\frac{ 3 }{ 4 }x+y=4\] Which is the wrong answer. Can some one help me find my error and get the right answer please?

OpenStudy (anonymous):

xray of a broken heart

OpenStudy (wolf1728):

Start with the formula b = y - mx b = 5 -.75*(-2) b = 5 + 1.5 b = 6.5 y = mx + 6.5 y = .75x + 6.5

OpenStudy (anonymous):

\[y-5=\frac{3}{4}(x+2)\]multiply out on the right \[y=5=\frac{3}{4}x+\frac{3}{2}\] add \(5\) to both sides \[y=\frac{3}{4}x+\frac{13}{2}\]

OpenStudy (anonymous):

i am not sure what "general form" means, but the above form is called "slope - intercept" form

OpenStudy (mertsj):

\[y-5=\frac{3}{4}(x+2)\] \[y-5=\frac{3}{4}x+\frac{3}{2}\] \[y=\frac{3}{4}x+\frac{13}{2}\] \[4y=3x+26\]\\[3x-4y=-26\]

OpenStudy (anonymous):

Ok thank you that is what I was missing was the step with the fractions.

OpenStudy (mertsj):

yw